# LLaVA-UHD v2: an MLLM Integrating High-Resolution Feature Pyramid via Hierarchical Window Transformer

### v1

@@ 4.3 Main Performance
LLaVA-UHD 58.6 70.0 63.8 55.4 56.3 62.2 56.7 45.6
w. ConvNext 59.7 68.2 62.7 55.6 61.8 63.5 61.8 44.0
w. DeConv. 61.7 71.2 64.2 57.4 61.8 67.8 63.4 46.3
w. Bilinear 62.0 72.0 64.5 57.8 62.2 67.6 63.7 46.5
w. JBU module 63.0 73.0 64.6 58.3 62.5 68.5 65.0 48.9
4.4 Analytical Study
We conduct analytical experiments on proposed modules to validate the effectiveness of LLaVA-UHD v2. Note that we freeze the CLIP-ViT during the supervised fine-tuning, which facilitates a clearer verification of the effectiveness of each module introduced in the Hiwin transformer.
Main module ablation. In Table 2, by replacing the low-resolution CLIP-ViT features with JBU-enhanced high-resolution ones, 2% average improvement can be seen, especially, on tasks depending on visual details like OCR-Bench (+9.5%) and RefCOCOs (+4.0%). Employing the Hiwin transformer to integrate the feature pyramid constructed by the JBU module further improves average accuracy by 1.5%, especially 5.0% on ultra-high resolution perception (HR-Bench), demonstrating rich visual granularity facilitates more precise language generation. Additionally, spatially consistent token organization leads to further performance improvements in visual-spatial understanding, with a 1.4% increase on RealWorldQA and a 0.7% improvement on RefCOCOs.
Figure 6: Qualitative comparison of proposed LLaVA-UHD v2 and advanced MLLMs, including LLaVA-Next, Mini-Gemini, and GPT-4V on high-resolution complex perception tasks, which require the integration of both fine-grained visual information and high-level semantic contexts.
Table 4: Comparison of different choice of grid sizes on performance and efficiency. “Pyramid” means the feature grids from different levels form a region-level feature pyramid, $e.g.$, [2$\times$3] for level-0, [4$\times$6] for level-1, [8$\times$12] for leval-2. “Fix” represents all feature maps are pooled into a 3$\times$3 feature grid. We measure the training period on 8$\times$×A100s, the latency on an A100 with a 1008$\times$672 image, and the GPU memory on 8$\times$×A100s with 1 image per GPU in supervised fine-tune phase.
Efficiency General Knowledge OCR & Chart
Pyramid 62.4 1.26 60.3 62.4 69.0 60.8 57.3 60.7 67.5 58.9
Fix [3$\times$3] 26.9 0.62 41.7 64.6 73.8 63.9 58.8 60.9 66.2 63.8
Selective 27.7 0.54 39.4 65.3 73.0 64.6 58.3 62.5 68.5 65.0
@@ 4.4 Analytical Study
JBU module demonstrates effectiveness on both MLLM tasks and vision tasks. We investigate the generalizability of different feature up-sampling methods ($i.e.$, bilinear interpolation, de-convolution [36], JBU module, and CNN [68]) across diverse tasks. (1) Results on MLLM tasks. Table. 3 clearly shows that feature pyramids, regardless of their construction method, can enhance performance across various MLLM tasks. Nonetheless, the feature pyramid constructed with our JBU module achieves an average performance gain of 1.0% over bilinear interpolation, indicating that the JBU module further enhances beneficial visual representations ($e.g.$, high-frequency visual features). (2) Results on visual tasks. Beyond the VQA task in MLLM, we further select some fundamental visual tasks, including semantic segmentation [58], optical character recognition [80], and fine-grained classification [45], to compare the effect of bilinear interpolation and JBU module. The settings of training and testing are detailed in supplement A.2. As shown in Fig. 5, the JBU module outperforms the bilinear interpolation on OCR (+4.0%), semantic segmentation (+4.7%) and fine-grained classification (+3.8%), demonstrating that the JBU module captures more visual details, enabling precise semantic discrimination.
The choice of grid sizes influences the performance of MLLMs. We explore the impact of RoI-align grid size on the efficiency and performance of LLaVA-UHD v2. Specifically, we RoI-align a region-level feature pyramid from the inverse feature pyramid within a window set, ensuring that higher-resolution feature maps retain finer-grained pooling grids, like [96]. However, this approach, rather than improving multi-scale feature integration, significantly degrades performance and inference efficiency, as demonstrated in Table 4. Compared to fixed grids, selecting a proper pooling grid (defined in Eq.5), offers better performance and efficiency, thanks to a more approximate aspect ratio to the original image.
@@ 4.5 Visualization Analysis
Case study. In Fig. 6, we visualize the performance of well-known MLLMs on high-resolution, complex perception tasks. This kind of task requires MLLMs to well fuse both visual details and high-level semantics to accurately identify fine-grained targets ($e.g.$, OCR, colors) during the procedure of complex semantic perception ($e.g.$, semantic relation and visual behavior). It is evident that LLaVA-UHD v2 correctly recognizes the tree planter in the newspaper photo and associates it with the name within the dense image caption (Case A). We also can see that LLaVA-UHD v2 captures the player who raises his hands and reads the “number 3” on his clothes (Case B). In contrast, LLaVA-Next overlooks the name information within dense texts (Case A) and hallucinates on the player number (Case B). Mini-Gemini fails to extract the true name (Case A) and also hallucinates (Case B). Additionally, GPT-4V shows limitations in referencing the information in the newspaper (Case A) and falsely recognizes “number 24” due to wrong fine-grained action perception (Case B). Appendix B.2.2 provides further cases for a more comprehensive comparison across various task scenarios.
Figure 7: Activation response of specific textual tokens to different visual feature levels. Red circles highlight the obvious difference between levels. (Best viewed in color and zoomed-in)
Semantic activation cross feature scale. In Fig. 7, we demonstrate the activation responses of specific textual prompts in the language model to the inverse feature pyramid. As shown, OCR-like textual tokens yield finer-grained and more accurate activations at higher feature levels, facilitating accurate scene text recognition (first row). For object-level semantics, higher feature levels enhance edge detail activations, enabling more precise semantic localization(second row). Collectively, the feature pyramid offers a more exhaustive set of visual semantics with rich granularity, effectively supporting nuanced language decoding.
@@ 5 Conclusion
The proposed hierarchical window transformer, central to our LLaVA-UHD v2, effectively addresses the limitations of conventional ViT-based MLLMs by capturing varying visual granularity essential for precise language generation. The Hiwin transformer adeptly constructs an inverse feature pyramid for enriched semantics representation, which is subsequently condensed into a compact set of visual tokens through our hierarchical window attention. This process enhances nuanced visual-linguistic alignment as well as facilitates efficient visual prompting for the LLM. LLaVA-UHD v2 has demonstrated substantial gains over the baseline method across a range of MLLM benchmarks, demonstrating its capability in MLLM tasks that demand both fine-grained and high-level semantics. Furthermore, the Hiwin transformer offers versatility, presenting potential adaptability across diverse ViT-based MLLM architectures.
@@ Appendix A Implementation Details
We provide details of the JBU module, experimental setting on visual tasks, and the dataset composition in the supervised fine-tuning phase of MLLM.
A.1 JBU module
Up-sampling kernel. As presented in Eq. 2 of the main paper, the kernel of the JBU module relies on two weights $D_{dist}$ and $D_{sim}$. Spatial distance decay $D_{dist}$ is defined as
where $\sigma_{sim}$ is a learnable temperature factor to modulate the distribution of similarity scores and $U=7$
Learnable down-sampler. We detail the implementation of the learnable down-sampler defined in Eq. 4 of the main paper. Compared to a simple convolutional layer with a stride of 2, we apply an attention-based down-sampler following  [26] in each feature level. Specifically, a 1$\times$1 convolution layer is first applied to the feature maps of $(l+1)$-th level to extract a saliency map, followed by combining it with a modified fully-connected layer to normalize the features in the local neighborhood $V$. We summarize the above operation as a network $f(\cdot)$ with trainable parameters $\Omega^{l+1}$. As a result, the feature pixel at the location of $[x,y]$ on down-sampled feature maps (of $l$-th level) is formally defined as
Multi-view hierarchical reconstruction loss. The objective of the JBU module is to construct high-resolution features by observing multiple different “views” of low-resolution features like NeRF [79] that generates 3D scene representation from multi-view 2D images. Given the input image $\mathcal{I}$ and $l$-th level of feature maps $\mathcal{F}^{l}$, we detail the Eq. 4 of the main paper as
$\mathcal{L}=\frac{1}{2|\mathcal{T}|}\sum_{l=1}^{2}\sum_{t\in\mathcal{T}}\bigg{(}\frac{1}{{{u}}^{2}}\lVert\mathcal{F}^{0}\left(t\left(\mathcal{I}\right)\right)-\mathrm{Down}\left(t\left(\mathcal{F}^{l}\right);\Omega^{l}\right)\rVert_{2}^{2}\\ +\log({u})\bigg{)},$
where $t$ is the image jitter operation sampled from a collection of transformation $\mathcal{T}$ such as pads, zooms, crops, flips, and $etc$. We experimentally set $|\mathcal{T}|=2$. The semantic uncertainty $u=\mathcal{N}(\mathcal{F}^{0}(t(\mathcal{I})))$, as introduced by [33], is parameterized by a linear network $\mathcal{N}$. This term evaluates the certainty of features for up-sampling operations, as such processes may risk losing critical semantic information, which is vital for preserving the integrity of the representation.
@@ A.1 JBU module
A.2 Experimental setting on visual tasks
After training the JBU module in Stage 0, we applied it on the top of the visual backbone to assess its effectiveness. Unless explicitly stated otherwise, both naive bilinear interpolation and the JBU module are employed to perform feature up-sampling on low-resolution (i.e., 24$\times$24) feature maps which are extracted from the second-to-last layer of CLIP-ViT [85].
[…]
Table 6: Comparison of different choices of feature level on performance and efficiency. “HS.”: hierarchical supervision. ESTVQA [101] is a VQA benchmark focusing on scene text recognition.
Efficiency Performance
Level Period(h) Memory(G) Average GQA SQA REC VQAC VQAT ESTVQA MMEP
0,2 27.7 41.9 63.4 63.9 69.5 71.5 60.5 66.5 40.6 71.0
0,1,2 28.0 41.9 63.7 63.8 70.2 71.8 60.5 66.9 40.8 72.1
0,1,2,3 45.6 53.0 63.8 64.4 69.3 72.6 60.7 66.4 41.6 71.4
0,1,2,3 (w/o HS.) 45.6 52.6 62.4 63.6 69.8 67.1 57.8 66.5 39.9 72.0
@@ A.2 Experimental setting on visual tasks
Figure 11: PCA visualization of the up-sampled features by JBU module on nature scene. With hierarchical supervision, the high-resolution features (8×8$\times$8 ×) could clearly depict object boundary and text appearance. (Best viewed in color and zoomed in)
Figure 12: PCA visualization of the up-sampled features by JBU module on OCR scene. With hierarchical supervision, the high-resolution features (8×8$\times$8 ×) could clearly depict object boundary and text appearance. (Best viewed in color and zoomed in)
@@ Appendix B Analysis
We analyze the behaviors of LLaVA-UHD v2 through qualitative and quantitative experiments.
@@ B.1 Quantitative experiment
Level choice. As shown in Table 6, the introduction of higher level (higher resolution) feature maps results in consistent enhancement of the average performance. However, incorporation of even higher resolution features, such as level-3 feature maps ($i.e.$, 8×8$\times$8 × resolution than level-0), yields marginal benefits while substantially increasing the training cost. If we abandon the hierarchical supervision strategy, performance suffers, which we attribute to the detailed degeneration issue discussed in Sec.3.2 of the main paper. In this case, the JBU module encounters significant challenges in effectively integrating high-frequency information into CLIP-ViT features.
@@ B.2 Qualitative experiment
As discussed in Sec.3.2 of the main paper, we performed a qualitative visualization to assess the impact of hierarchical supervision, presented in Fig.11 and Fig.12, Note that all the high-resolution features are 8×8$\times$8 × resolution than CLIP-ViT features. While bilinear interpolation increases the nominal resolution of features, it fails to enhance the fidelity of image detail representation. In comparison, a naive JBU module captures finer details but retains a degree of blurriness. With hierarchical supervision, the degradation of high-resolution features is markedly reduced, resulting in a more accurate and refined representation of visual details.
@@
We add more cases to analyze the behavior of our LLaVA-UHD v2. The capabilities are summarized as four aspects as follows.
@@
It is evident that LLaVA-UHD v2 and GPT-4V accurately identify the beginning time of the television program ‘H20 X5 Mop’ (case A), the performing date of the show ‘The 3 Mile’ (case B), the duration of whole workouts (case C), and the prize of stoli (case D), indicating highly robust fine-grained perception capabilities provided by our visual pyramid representation. In comparison, other models either fail to precisely locate the target (LLaVA-Next) or cannot distinguish the target from similar adjacent objects, limited in accurately completing dense OCR tasks (Mini-Gemini).
@@
Cases C indicate that LLaVA-UHD v2 accurately identified the small green light, and the tiny number of duration time associated with green light, demonstrating that the introduction of high-frequency information in hierarchical features can handle small, blurry targets effectively. In contrast, other models, can’t find the small green light, or fail to accurately perform OCR tasks due to the text being too small or blurry (e.g., GPT-4V, LLaVA-Next, Mini-Gemini).

### v2

@@ 4.4 Analytical Study
LLaVA-UHD 59.9 70.0 63.8 55.4 56.3 62.2 56.7 45.6
w/ ConvNext 59.7 68.2 62.7 55.6 61.8 63.5 61.8 44.0
w/ DeConv. 61.7 71.2 64.2 57.4 61.8 67.8 63.4 46.3
w/ Bilinear 62.0 72.0 64.5 57.8 62.2 67.6 63.7 46.5
w/ VDIM 63.0 73.0 64.6 58.3 62.5 68.5 65.0 48.9
@@
ISP demonstrates effectiveness on MLLM tasks compared to traditional feature pyramid. In Table. 3, it is evident that feature pyramids, regardless of their construction method, could enhance performance across various MLLM tasks. Nonetheless, the ISP constructed by VDIM achieves an average performance gain of 1.0% over bilinear interpolation, indicating that the ISP further enhances beneficial visual representations ($e.g.$, high-frequency visual details).
Figure 5: Comparison of performance. (a) Performance of using different projectors on compressing ISP. Hiwin attention exhibits a significant advantage. (b) Performance of our model equipped with different LLMs.
Hierarchical window attention works on compressing multi-scale feature pyramid. As shown in Fig. 5(a), Hiwin attention condenses visual tokens by using learnable queries on feature pyramids with arbitrary native resolutions, resulting in superior performance on fine-grained MLLM tasks (e.g., DocVQA and TextVQA). In contrast, directly down-sampling high-resolution features via interpolation before MLP projection [67] loses critical visual details and thereby degrades performance. Moreover, the Perceive Resampler [6, 110] performs notably worse due to its lack of spatial prior constraints on each query, which adversely affects training convergence, as evidenced by its significantly higher pre-training loss compared to Hiwin attention (0.6110 vs. 0.4368).
Hiwin transformer keeps superior when transferring to other LLMs. In Fig. 5(b), the proposed Hiwin transformer consistently improves performance across Vicuna-13B, and Qwen2-7B, demonstrating its strong generalization capability to different model scales and architectures. Especially in tasks that require visual details, such as DocVQA and HR-Bench, the improvement is even more pronounced.
@@ 4.5 Visualization Analysis
Figure 6: Qualitative comparison of proposed LLaVA-UHD v2 and advanced MLLMs, including LLaVA-Next, Mini-Gemini, and GPT-4V. Our method outperforms its counterparts, by providing both fine-grained visual information and high-level semantic contexts for the high-resolution complex perception tasks
Figure 7: Activation response of specific textual tokens to different visual feature levels, exhibiting complement to each other. Red circles highlight the obvious difference between levels (Best viewed in color and zoomed-in).
Case study. In Fig. 6, we visualize the performance of well-known MLLMs on high-resolution, complex perception tasks. This kind of task requires MLLMs to well fuse both visual details and high-level semantics to accurately identify fine-grained targets ($e.g.$, OCR, colors) during the procedure of complex semantic perception ($e.g.$, semantic relation and visual behavior). It is evident that LLaVA-UHD v2 correctly recognizes the tree planter in the newspaper photo and associates it with the name within the dense image caption (Case A). We also can see that LLaVA-UHD v2 captures the player who raises his hands and reads the “number 3” on his clothes (Case B). In contrast, LLaVA-Next overlooks the name information within dense texts (Case A) and hallucinates on the player number (Case B). Mini-Gemini fails to extract the true name (Case A) and also hallucinates (Case B). Additionally, GPT-4V shows limitations in referencing the information in the newspaper (Case A) and falsely recognizes “number 24” due to wrong fine-grained action perception (Case B).
Semantic activation cross semantic scale. In Fig. 7, we demonstrate the activation responses of specific textual prompts in the language model to the inverse semantic pyramid. As shown, OCR-like textual tokens yield finer-grained and more accurate activations at higher feature levels, facilitating accurate scene text recognition (first row). For object-level semantics, higher feature levels enhance edge detail activations, enabling more precise semantic localization (second row). Collectively, the semantic pyramid offers a more exhaustive set of visual semantics with rich granularity, effectively supporting nuanced language decoding.
@@ 5 Conclusion
Our proposed hierarchical window transformer, which is the core of LLaVA-UHD v2, effectively addresses the limitations of conventional ViT-based MLLMs by capturing varying visual granularity essential for precise language generation. The Hiwin transformer adeptly constructs an inverse semantic pyramid for enriched multi-modal representation, which is then condensed into a compact set of visual tokens. This process enhances nuanced visual-linguistic alignment as well as facilitates efficient visual prompting for the LLM. LLaVA-UHD v2 shows substantial gains over the baseline method across a range of MLLM benchmarks, demonstrating its capacity in MLLM tasks that demand both low-level details and high-level semantics. Furthermore, the Hiwin transformer offers versatility, presenting potential adaptability across diverse ViT-based MLLM architectures.
@@ Appendix A Implementation Details
A.1 Experimental setting on visual tasks
Results on visual tasks. Beyond the VQA task in MLLM, we further evaluate some fundamental visual tasks, including semantic segmentation [62], optical character recognition [86], and fine-grained classification [49], to compare the effect of bilinear interpolation and VDIM. In Fig. 8, the VDIM outperforms the bilinear interpolation on OCR (+4.0%), semantic segmentation (+4.7%) and fine-grained classification (+3.8%), which demonstrate more visual details are encoded for precise semantic discrimination. The implementation details for each visual task are provided below. Unless explicitly stated, both naive bilinear interpolation and the VDIM are employed to perform feature up-sampling on low-resolution (i.e., 24$\times$24) feature maps which are extracted from the second-to-last layer of CLIP-ViT [91].
Optical character recognition. We follow the experimental setting of [10], training on MJSynth [43] dataset and evaluating on IIIT5K [87] dataset. Specifically, we first up-sample the low-resolution feature maps to high-resolution ones ($i.e.$, 48$\times$48) by using bilinear interpolation and a pre-trained VDIM. The resulting feature maps are then fed to a sequence encoder and a CTC head [33] to predict the class labels of characters in the images.
[…]
Linear probing semantic segmentation. We follow the experimental setting of previous research [29, 5, 38]. To be specific, on the COCOStuff dataset [62], we train a linear projection upon a frozen CLIP-ViT to directly predict the class label of each pixel.
[…]
Figure 8: Performance on different visual tasks with VDIM and vanilla bilinear interpolation. “OCR" denotes the optical character recognition, “Seg" the Linear probing semantic segmentation, and “Cls" the fine-grained classification on SUB-200.
Table 4: Comparison of different choices of feature level on performance and efficiency. ESTVQA [106] is a VQA benchmark focusing on scene text recognition.
Efficiency Performance
Level
Period(h)
Memory(G)
Average
GQA
SQA
REC
VQAC
VQAT
ESTVQA
MMEP
0,2 27.7 41.9 63.4 63.9 69.5 71.5 60.5 66.5 40.6 71.0
0,1,2 28.0 41.9 63.7 63.8 70.2 71.8 60.5 66.9 40.8 72.1
0,1,2,3 45.6 53.0 63.8 64.4 69.3 72.6 60.7 66.4 41.6 71.4
0,1,2,3 (w/o MRL) 45.6 52.6 62.4 63.6 69.8 67.1 57.8 66.5 39.9 72.0
@@ A.1 Experimental setting on visual tasks
Table 6: Comparison of different choices of grid sizes on performance and efficiency. “Pyramid" means the feature grids from different levels form a region-level feature pyramid, $e.g.$, [2$\times$3] for level-0, [4$\times$6] for level-1, [8$\times$12] for level-2. “Fix" represents all feature maps are pooled into a 3$\times$3 feature grid. “Selective" represents adaptively selecting a grid size that is closest to the resolution. We measure the training period on 8$\times$×A100s, the latency on an A100 with a 1008$\times$672 image, and the GPU memory on 8$\times$×A100s with 1 image per GPU in supervised fine-tun ing phase.
Efficiency General Knowledge OCR & Chart
Pyramid 62.4 1.26 60.3 62.4 69.0 60.8 57.3 60.7 67.5 58.9
Fix [3$\times$3] 26.9 0.62 41.7 64.6 73.8 63.9 58.8 60.9 66.2 63.8
Selective 27.7 0.54 39.4 65.3 73.0 64.6 58.3 62.5 68.5 65.0
@@
Figure 12: PCA visualization of the up-sampled features by VDIM on nature scene. With VIDM, the high-resolution features could clearly depict object boundaries and text appearance. (Best viewed in color and zoomed in)
Figure 13: PCA visualization of the up-sampled features by VDIM on OCR scene. With VDIM, the high-resolution features could clearly depict object boundaries and text appearance. (Best viewed in color and zoomed in)
A.2 Visual detail injection module (VDIM)
In this setction, we detail the implementation of the proposed VDIM. Specifically, each pixel value of the up-sampled feature maps is extracted as
where $[x,y]$ denotes the coordinate index of feature maps, $U$ the region of convolution neighborhood, $D_{dist}$ is a decay factor based on the distance between $[x,y]\ \text{and}\ [x^{\prime},y^{\prime}]$, and $D_{sim}$ is a similarity weight based on the attention of each image pixel using a fully connected layer $\Theta^{l+1}$. By iteratively performing Eq. 6, we progressively up-sample the feature maps and finally construct the ISP $\{\mathcal{F}^{0},\mathcal{F}^{1},\mathcal{F}^{2}\}$.
Up-sampling kernel. Following [37], the kernel weight $\Theta^{l+1}$ in Eq.6 of the VDIM relies on two weights $D_{dist}$ and $D_{sim}$. Spatial distance decay $D_{dist}$ is defined as
where $\sigma_{sim}$ is a learnable temperature factor to modulate the distribution of similarity scores and $U=7$.
Learnable down-sampler. We detail the implementation of the learnable down-sampler defined in Eq.2 of the main paper. Compared to a simple convolutional layer with a stride of 2, we apply an attention down-sampler following  [29] at each feature level.
Specifically, a 1$\times$1 convolution layer is first carried out on the feature maps of $(l+1)$-th level to extract a saliency map, followed by combining it with a modified fully-connected layer to normalize the features in the local neighborhood $V$. We summarize the above operation as a network $f(\cdot)$ with trainable parameters $\Omega^{l+1}$. As a result, the feature pixel at the location of $[x,y]$ on down-sampled feature maps (of $l$-th level) is formally defined as
@@ B.1 Quantitative experiment
Level choice. As shown in Table 4, the introduction of higher level (higher resolution) feature maps results in consistent enhancement of the average performance. However, incorporation of even higher resolution features, such as level-3 feature maps ($i.e.$, 8×8$\times$8 × resolution than level-0), yields marginal benefits while substantially increasing the training cost.
When the multi-level reconstruction (MLR) supervision is replaced by the last-level supervision, performance deteriorates. In such a scenario, the VDIM faces substantial challenges in effectively incorporating high-frequency information into CLIP-ViT features.
The choice of RoI-align grid sizes We explore the impact of RoI-align grid size on the efficiency and performance of LLaVA-UHD v2. Specifically, we RoI-align a region-level feature pyramid from the inverse feature pyramid within a window set, ensuring that higher-resolution feature maps retain finer-grained pooling grids, like [102]. However, this approach, rather than improving multi-scale feature integration, significantly degrades performance and inference efficiency, as demonstrated in Table 6. Compared to fixed grid, selecting a proper pooling grid (defined in Eq.3 of main paper) showcases better performance and efficiency, because of a more approximate aspect ratio to the native image.
@@ B.2 Qualitative experiment
We performed a qualitative visualization of our ISP, presented in Fig.12 and Fig.13, Note that all the high-resolution features shown are the highest-level of ISP enhanced from CILP-ViT features. While bilinear interpolation increases the nominal resolution of features, it fails to enhance the fidelity of image detail representation. Using the FeatUp [29] for feature up-sampling, the resulting features capture finer details but retain a degree of blurriness. In contrast, with the proposed VIDM, the high-resolution features clearly depict object boundary and text appearance, demonstrating accurate and refined representation of visual details.
@@
We add more cases to analyze the behavior of our LLaVA-UHD v2. The capabilities are summarized as three aspects as follows.
@@
It is evident that LLaVA-UHD v2 and GPT-4V accurately identify the beginning time of the television program ‘H20 X5 Mop’ (case A), the performance date of the show ‘The 3 Mile’ (case B), the duration of whole workouts (case C), and the prize of Stoli (case D), indicating highly robust fine-grained perception capabilities provided by our visual pyramid representation. In comparison, other models either fail to precisely locate the target (LLaVA-Next) or cannot distinguish the target from similar adjacent objects, limited in accurately completing dense OCR tasks (Mini-Gemini).
@@
Case C indicates that LLaVA-UHD v2 accurately identified the small green light, and the tiny number of duration time associated with green light, demonstrating that the introduction of high-frequency information in hierarchical features can handle small, blurry targets effectively. In contrast, other models can not find the small green light, or fail to accurately perform OCR tasks due to the text being too small or blurry (e.g., GPT-4V, LLaVA-Next, Mini-Gemini).